Find the solutions of the equation if 0≤𝑡<2𝜋

tan^2 𝑡−1=0



1
Expert's answer
2021-11-30T04:53:42-0500
"\tan^2t-1=0"

"(\tan t-1)(\tan t+1)=0"

"\tan t=\pm1"

"t=\dfrac{\pi}{4}+\dfrac{\pi n}{2}, n\in \Z"

If we consider "0\leq t<2\pi"

"t_1=\dfrac{\pi}{4}, t_2=\dfrac{3\pi}{4}, t_3=\dfrac{5\pi}{4}, t_4=\dfrac{7\pi}{4}"

"\{\dfrac{\pi}{4}, \dfrac{3\pi}{4}, \dfrac{5\pi}{4}, \dfrac{7\pi}{4}\}"